OpenStudy (anonymous):

What is the approximate volume of the cylinder? Use 3.14 to approximate . A. 160 ft3 B. 214 ft3 C. 427 ft3 D. 2722 ft3 I think the correct answer would be C The Base of the circle is 17 height is 3

OpenStudy (anonymous):

\[V=\pi \times r ^{2}\times h\] By base, do you mean radius?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

\[V=\pi \times 17^{2} \times 3\]\[17^{2}=289\]\[\pi \times 289 \times 3=2722.38\]

OpenStudy (anonymous):

Thanks I understand how to do it now :D

OpenStudy (anonymous):

That's good :)